Pharmacy for the homeless is going to purchase a new drug storage unit for $16,000. A bank will lend the organization the case at an annual interest rate of 12%. Pharmacy will have to repay the loan in equal installments over two years. How much will they have to pay each month?
1
Expert's answer
2013-06-17T08:48:25-0400
The full sum to repay in two years will be:
FV = PV*(1+r)^2 = 16,000*1.12^2 = $23,030, FV - future value, PV - present
value, r - interest rate.
So, the monthly installment amount will be:
MI = FV/24 = 23,030/24 = $960 they have to pay each month